Instructions: solve each system by the eliminsation method.
What is the solution for x-y=4
x-y=-3

Answer by funmath(2933)

Instructions: solve each system by the eliminsation method.
What is the solution for x-y=4
x-y=-3
Don't freak out, but this one has no solution. If you try to eliminate either variable both go and you get 0=7 or 0=-7. Graphically, these lines are parallel and will never intersect.
Multiply the top equation by -1 and add them together.
-1(x-y)=-1(4)--->-x+y=-4
:
-x+y=-4
x-y=-3
_______
0+0=-7
0=-7 No solution.
Happy Calculating!!!